# Triplet scalar flavored leptogenesis with spontaneous CP violation

### v1

@@ 1 Introduction
The coupling of the triplet scalars to the leptons and the SM Higgs scalar leads to two different decay channels of the triplet scalar and the lepton number is violated only if both types of decays co-exist. The interplay of the strength of the respective branching ratios, $B^{L}$ and $B^{\phi}$ plays an important role in successful baryogenesis through leptogenesis, where $B^{L}(B^{\phi})$ corresponds to the branching ratios of the triplet scalar to the leptons (SM Higgs scalar). The amount of CP asymmetry, $\epsilon$, and the efficiency of leptogenesis depend on $B^{L}$ and $B^{\phi}$ in a different manner. The efficiency is maximal when $B^{L}\gg B^{\phi}$ or $B^{\phi}\gg B^{L}$, whereas this condition corresponds to suppressed CP violation (as $\epsilon\propto\sqrt{B^{L}B^{\phi}}$). Solving a full set of Boltzmann equations, in a model-independent way, it was shown in Ref. [31] that even for extremely hierarchical branching ratios, it is possible to achieve the required BAU. In this work, we focus on leptogenesis from the decay of triplet scalars with hierarchical branching ratios and CP violation originating from the VEV of a scalar singlet, embedded in the SM along with with two triplet scalars. Also to make the model more predictive we impose $A_{4}$ symmetry at the Lagrangian level. Since in our case, adequate baryon asymmetry is obtained for the triplet scalar mass around $\sim 10^{10}$ GeV, flavor effects become important in this regime. Also, considering hierarchical branching ratios of triplet scalar decays, washout effects play an important role in generating adequate lepton/baryon asymmetry. With the improved computation of the kinetic evolution of flavored lepton asymmetries, we use Density Matrix Equations (DME) for the evaluation of final lepton asymmetries. We also study, with all possible mass hierarchies among the triplet scalars, the flavor structure of the model predicts different scales of leptogenesis.
The paper is organized as follows. The model is introduced in section (2). Section (3) is devoted to pointing out the important features of flavored CP asymmetry and DME for the analysis. Finally, the results are presented in section (4) with the summary. The conclusion is given in section (5). Three appendices (A), (B), (C) are added containing some of the detailed calculations.
Field
$L$
$e_{R},\mu_{R},\tau_{R}$
$\Delta_{1}$
$\Delta_{2}$
$\phi$
$S$
$\Phi$
$\psi$
$SU(2)_{L}$
$2$
$1$
$3$
$3$
$2$
$1$
$1$
$1$
$U(1)_{Y}$
$-1/2$
$-1$
$1$
$1$
$1/2$
00
00
00
$A_{4}$
$3$
$1,1^{\prime\prime},1^{\prime}$
$1$
$1$
$1$
$1$
$3$
$3$
$Z_{4}$
$i$
$-i$
$1$
$-1$
$i$
$-1$
$i$
$1$
@@ 2 Model Description
$V_{0}\simeq-\frac{m^{4}_{S}}{4(\lambda_{S}-2\lambda^{\prime}_{S})}$ (2.11)
@@
where $m^{(1)}_{\nu}$ and $m^{(2)}_{\nu}$ are the contributions coming from the triplet scalars $\Delta_{1}$ and $\Delta_{2}$, respectively. Here, $u_{a}$ is the VEV of the neutral components of the triplet scalars $\Delta_{a}$, and it is given by
@@
TBM mixing matrix has gathered considerable attention because of its proximity with the observation of neutrino masses and mixing. However the absence of Dirac-type CP violating phase asks for possible modifications of the model, which we discuss in the next subsection. Considering the two cases of neutrino mass hierarchy, we can have two sets of constraints on $z_{1}$ and $z_{2}$. From Eq.(2.14), for normal hierarchy (NH), i.e. $m_{3}>m_{2}>m_{1}$, we obtain
@@ 2.2 Parameter space study
[…]
$\sin^{2}\theta_{12}$
$0.303^{+0.012}_{-0.012}$
$0.270\rightarrow 0.341$
$\sin^{2}\theta_{23}$
$0.451^{+0.019}_{-0.016}$
$0.408\rightarrow 0.603$
$\sin^{2}\theta_{13}$
$0.02225^{+0.00056}_{-0.00059}$
$0.02052\rightarrow 0.02398$
$\delta/^{\circ}$
$232^{+36}_{-26}$
$144\rightarrow 350$
$\frac{\Delta m^{2}_{21}}{10^{-5}{\rm eV}^{2}}$
$7.41^{+0.21}_{-0.20}$
$6.82\rightarrow 8.03$
$\frac{\Delta m^{2}_{3l}}{10^{-3}{\rm eV}^{2}}$
$+2.507^{+0.026}_{-0.027}$
$+2.427\rightarrow+2.590$
@@
where the phase $\gamma_{1}$ is given in Eq. (2.19). Through the phase, $\gamma_{1}$, it is possible to draw a link between the effective mass and its dependence on the phase $\beta$, which appears in the expressions of neutrino masses as well as in the high-scale CP violation, as it will be elaborated later on. In Fig. (0(b)) we show the dependence of $|m_{ee}|$ on the phase $\beta$. The plot shows that the allowed effective mass in turn has imposed a bound on the phase $\beta\leq 270^{\circ}$. Apart from the bound on the effective neutrino mass $|m_{ee}|\lesssim 0.15$ eV, the bound on the total neutrino mass $\Sigma=m_{1}+m_{2}+m_{3}<0.12$ eV [6] puts a more stringent restriction on the parameter space of phase $\beta$ as $\sim[120^{\circ},240^{\circ}]$. We have used this restricted parameter space of the phase $\beta$ to calculate the flavored CP asymmetries in section (4).
@@ 4.1 Formulation of flavored CP asymmetries
Within the energy range $[10^{9},10^{12}]$ GeV, the flavored CP asymmetries can be expressed as $\epsilon^{\alpha\beta}_{a}(\alpha,\beta=e,\mu)$ and $\epsilon^{\tau\tau}_{a}$, where $a=1,2$ is associated with triplet $\Delta_{1,2}$. Using equations (4.3), (4.4) and (4.6) we calculate, the flavored CP asymmetries. The specific expressions are given in equations (C.8 - C.11) and (C.12 - C.15) in Appendix (C). The dependence of $\beta$ on the CP asymmetries are plotted in the figures (8) and (9) for different values $B^{L}_{1(2)}=0.5,0.05,0.005,0.0005$. The triplet VEV $|u_{1}|$ ($|u_{2}|$) are taken in the range $[0.1-100]$ eV. To estimate the CP asymmetry from the model, we have chosen the small perturbation parameters from the allowed parameter space shown in Fig.(0(a)), as $\varepsilon_{1}\sim-0.125$ and $\varepsilon_{2}\sim 0.075$. We have varied the phase $\beta$ within the range $[120^{\circ},240^{\circ}]$ so that it also remains consistent with the latest cosmological bound on the sum of the neutrino masses and the effective neutrino mass $|m_{ee}|\lesssim 0.15$ eV. The phase $\sigma_{1}$ is taken in such a way that it satisfies the relations given in Eq.(2.19). The viable values of the parameters $z_{1}$ and $z_{2}$ are calculated as functions of two mass squared differences and phase $\beta$, from the expressions equations (2.22) and (2.23) by making sure that it is consistent with the bound on total neutrino mass $\Sigma=m_{1}+m_{2}+m_{3}<0.12$ eV through the expression in Eq.(2.24). It is observed that CP asymmetries of the order $\mathcal{O}(10^{-7})$ can be produced for triplet masses $\sim 10^{10}$ GeV. The figures (8) and (9) are generated by taking $M_{1}$ ($M_{2}$) $=2\times 10^{10}$ GeV and taking the mass squared differences in $3\sigma$ range. The scale is suitable for studying two-flavor leptogenesis which we take up in the upcoming sections.
@@ 4.2 Baryogenesis from $\Delta_{1}$ decay
$M_{1}({\rm GeV})$
$B^{L}_{1}$
$\beta(^{\circ})$
$\epsilon^{ee}_{1}/10^{-8}$
$\epsilon^{e\mu}_{1}/10^{-9}$
$\epsilon^{\mu\mu}_{1}/10^{-9}$
$\epsilon^{\tau\tau}_{1}/10^{-9}$
$|\eta_{B}|/10^{-10}$
$0.48$
$178$
$0.70$
$0.84$
$1.01$
$-1.68$
$0.04$
$3\times 10^{10}$
$0.06$
$175$
$2.06$
$2.46$
$2.94$
$-4.91$
$0.29$
$0.005$
$165$
$2.04$
$2.43$
$2.91$
$-4.86$
$0.73$
$0.0005$
$126$
$2.06$
$2.46$
$2.94$
$-4.91$
$1.74$
$0.34$
$178$
$2.79$
$3.32$
$3.98$
$-6.63$
$0.30$
$4\times 10^{10}$
$0.05$
$175$
$2.82$
$3.36$
$4.03$
$-6.71$
$0.44$
$0.005$
$164$
$2.82$
$3.36$
$4.03$
$-6.71$
$1.07$
$0.0006$
$126$
$2.83$
$3.37$
$4.04$
$-6.74$
$2.50$
@@
$M_{1}({\rm GeV})$
$B^{L}_{1}$
$|u_{1}|(\rm eV)$
$\epsilon^{ee}_{1}/10^{-7}$
$\epsilon^{e\mu}_{1}/10^{-8}$
$\epsilon^{\mu\mu}_{1}/10^{-8}$
$\epsilon^{\tau\tau}_{1}/10^{-8}$
$|\eta_{B}|/10^{-10}$
$3\times 10^{10}$ Fig. (4)
$0.58$
$4.80$
$4.47$
$5.32$
$6.38$
$-1.06$
$5.29$
@@ 4.3 Baryogenesis from $\Delta_{2}$ decay
$M_{2}({\rm GeV})$
$B^{L}_{2}$
$\beta(^{\circ})$
$\epsilon^{ee}_{2}/10^{-8}$
$\epsilon^{e\mu}_{2}/10^{-9}$
$\epsilon^{\mu\mu}_{2}/10^{-9}$
$\epsilon^{\tau\tau}_{2}/10^{-7}$
$|\eta_{B}|/10^{-10}$
$0.59$
$176$
$-0.66$
$-0.79$
$-0.95$
$0.75$
$0.04$
$1\times 10^{10}$
$0.06$
$172$
$-0.67$
$-0.80$
$-0.96$
$0.76$
$0.07$
$0.005$
$157$
$-0.67$
$-0.80$
$-0.96$
$0.75$
$0.86$
$0.0006$
$126$
$-0.73$
$-0.87$
$-1.05$
$0.75$
$3.26$
$0.53$
$177$
$-1.29$
$-1.53$
$-1.84$
$1.45$
$0.01$
$2\times 10^{10}$
$0.05$
$176$
$-1.42$
$-1.68$
$-2.02$
$1.59$
$0.03$
$0.006$
$156$
$-1.42$
$-1.69$
$-2.03$
$1.58$
$1.66$
$0.0005$
$126$
$-1.42$
$-1.69$
$-2.03$
$1.45$
$7.51$
@@
$M_{2}({\rm GeV})$
$B^{L}_{2}$
$|u_{2}|(\rm eV)$
$\epsilon^{ee}_{2}/10^{-8}$
$\epsilon^{e\mu}_{2}/10^{-9}$
$\epsilon^{\mu\mu}_{2}/10^{-9}$
$\epsilon^{\tau\tau}_{2}/10^{-7}$
$|\eta_{B}|/10^{-10}$
$1.4\times 10^{10}$ Fig. (5)
$0.0006$
$48.1$
$-1.00$
$-1.20$
$-1.44$
$1.04$
$4.92$

### v2

@@ 1 Introduction
The coupling of the triplet scalars to the leptons and the SM Higgs scalar leads to two different decay channels of the triplet scalar and the lepton number is violated only if both types of decays co-exist. The interplay of the strength of the respective branching ratios, $B^{L}$ and $B^{\phi}$ plays an important role in successful baryogenesis through leptogenesis, where $B^{L}(B^{\phi})$ corresponds to the branching ratios of the triplet scalar to the leptons (SM Higgs scalar). The amount of CP asymmetry, $\epsilon$, and the efficiency of leptogenesis depend on $B^{L}$ and $B^{\phi}$ in a different manner. The efficiency is maximal when $B^{L}\gg B^{\phi}$ or $B^{\phi}\gg B^{L}$, whereas this condition corresponds to suppressed CP violation (as $\epsilon\propto\sqrt{B^{L}B^{\phi}}$). Solving a full set of Boltzmann equations, in a model-independent way, it was shown in Ref. [31] that even for extremely hierarchical branching ratios, it is possible to achieve the required BAU. In this work, we focus on leptogenesis from the decay of triplet scalars with hierarchical branching ratios and CP violation originating from the VEV of a scalar singlet, embedded in the SM along with two triplet scalars. Also to make the model more predictive we impose $A_{4}$ symmetry at the Lagrangian level. Since in our case, adequate baryon asymmetry is obtained for the triplet scalar mass around $\sim 10^{10}$ GeV, flavor effects become important in this regime. Also, considering hierarchical branching ratios of triplet scalar decays, washout effects play an important role in generating adequate lepton/baryon asymmetry. With the improved computation of the kinetic evolution of flavored lepton asymmetries, we use Density Matrix Equations (DME) for the evaluation of final lepton asymmetries. We also study, with all possible mass hierarchies among the triplet scalars, the flavor structure of the model predicts different scales of leptogenesis.
The paper is organized as follows. The model is introduced in section (2). Section (3) is devoted to point out the important features of flavored CP asymmetry and DME for the analysis. Finally, the results are presented in section (4) with the summary. The conclusion is given in section (5). Three appendices (A), (B), (C) are added containing some of the detailed calculations.
Field $L$ $e_{R},\mu_{R},\tau_{R}$ $\Delta_{1}$ $\Delta_{2}$ $\phi$ $S$ $\Phi$ $\psi$
$SU(2)_{L}$ $2$ $1$ $3$ $3$ $2$ $1$ $1$ $1$
$U(1)_{Y}$ $-1/2$ $-1$ $1$ $1$ $1/2$ 00 00 00
$A_{4}$ $3$ $1,1^{\prime\prime},1^{\prime}$ $1$ $1$ $1$ $1$ $3$ $3$
$Z_{4}$ $i$ $-i$ $1$ $-1$ $i$ $-1$ $i$ $1$
@@ 2 Model Description
$V_{0}\simeq-\frac{m^{4}_{S}}{4(\lambda_{S}-2\lambda^{\prime}_{S})},$ (2.11)
@@
One of the constraints on models with spontaneous CP violation comes from the cosmological scenario of domain wall problem [54]. A review of topological defects like cosmic strings, domain walls, and monopoles arising from cosmological phase transition in the early universe can be found in Ref. [53]. The spontaneous breaking of a global discrete symmetry leads to the formation of domain walls [50, 51]. The energy density of the domain walls redshifts slower than those of matter and radiation. Long-lived domain walls conflict with standard big bang cosmology. There are several solutions to make them unstable. One of the solutions is assuming gravity breaks global discrete symmetries explicitly, which lifts the degeneracy of the two vacua, there will be no domain wall problem [52, 49]. Lifting the degeneracy of the two vacua by an amount $\epsilon$ generates a pressure difference between the two sides of the walls with a tendency to push the wall into the false vacuum. By making a careful analysis of the dynamics of evolution with domain walls in a radiation-dominated universe, it can be shown that the domain walls are pushed to false vacuum if
$\epsilon\geq\frac{v_{\rm dis}^{6}}{M^{2}_{\rm Pl}},$ (2.12)
where $v_{\rm dis}$ is the scale of discrete symmetry breaking and $M_{\rm Pl}\sim 10^{19}$ GeV. In our case, $v_{\rm dis}$ can be replaced by the SCPV scale, $v_{S}$. The authors of Ref. [52] have considered generating such gravity-induced terms and their effects in destabilizing the domain walls. For a theory with generic neutral scalar field $\phi$ the Planck-scale suppressed/gravity-induced operators can be written as
$\Delta V=\frac{c_{5}}{M_{\rm Pl}}\phi^{5}+\frac{c_{6}}{M^{2}_{\rm Pl}}\phi^{6}+{\rm Higher~{}order~{}terms}.$ (2.13)
In general, the coefficients $c_{n}$ are of order $1$. The terms on the right-hand side of Eq. (2.13) are generic. In a specific model such as the one studied in this paper, the effective higher degree potential induced by gravity has to be constructed carefully keeping gauge invariance intact. In the model under consideration, the Lagrangian has been written keeping $A_{4}\times Z_{4}$, CP and gauge invariance intact. Also, we have assumed that the singlet scalar $S$ is very heavy and decouples from the theory at an energy scale much higher than the electroweak and Higgs-triplet scales. In Appendix (A), the minimization of the scalar potential is done by taking only the complex vev of $S$ into account. Keeping these in mind, the first higher-dimensional operator that can break CP symmetry explicitly can be written in terms of the complex gauge singlet as
$\Delta V=\frac{c_{6}}{M^{2}_{\rm Pl}}(SS^{*})^{3}+{\rm H.c.}+{\rm Higher~{}order~{}terms}.$ (2.14)
Since gravity is expected to break discrete symmetry explicitly, CP symmetry breaking can be ensured by taking $c_{6}$ as complex. However, the successful removal of domain walls requires $c_{6}\gtrsim 1$ as per Eq.(2.12). So careful study of the detailed consequences of gravity is important in low energy issues such as CP violation. Nevertheless, gravity can break the $Z_{4}$ discrete symmetry ($S$ is charged under $Z_{4}$). Thus, in this case, it is possible to construct a dimension-five operator of the form
$\Delta V=\frac{c_{5}}{M_{\rm Pl}}\mathcal{O}+{\rm H.c.}+{\rm Higher~{}order~{}terms},$ (2.15)
where the operator $\mathcal{O}$ can be of the form $S^{5},(SS^{*})S^{3}$, and $(SS^{*})^{2}S$. CP symmetry is explicitly broken with the dimension-five operator with $c_{5}$ as complex. However by comparing the above contribution as per Eq.(2.12), the relation $c_{5}\gg\frac{v_{S}}{M_{\rm Pl}}$ has to hold. This ensures the dimension-five operator is much bigger than the amount $\frac{v_{S}^{6}}{M^{2}_{\rm Pl}}$. In our case the scale of leptogenesis is $>10^{10}$ GeV. So taking the $v_{S}$, the SCPV scale above the leptogenesis scale, the values of $c_{5}$ are bounded below as
$c_{5}>\begin{cases}10^{-7},&v_{S}\sim 10^{12},\\ 10^{-6},&v_{S}\sim 10^{13},\\ 10^{-5},&v_{S}\sim 10^{14}.\end{cases}$ (2.16)
The above bound is easily satisfied as $c_{5}$ is of order $1$. This makes the domain wall network disappear early enough for an acceptable leptogenesis scenario.
There are several other ways to deal with the domain wall problem. The domain walls can become unstable by the existence of a non-zero $\theta_{\rm QCD}$ through non-perturbative effects, that break the CP explicitly and lift the degeneracy of the two CP conjugate vacua [55]. Another idea in the same direction is to invoke inflation after the domain walls are produced [56]. In Ref. [57], the authors study the possibility of spontaneous discrete symmetry breaking due to a gauge singlet field whose initial value is fixed during inflation, so that the resulting discrete symmetry-breaking phase is the same over the whole observable universe. However, the baryon asymmetry has to be generated after inflation to avoid diluting it to an insignificant level. In most of the models on leptogenesis, the reheat temperature required is $10^{12}$ GeV (in our case the scale of leptogenesis is $\mathcal{O}(10^{10})$ GeV). So it is hard to make domain walls unstable unless they are associated with a spontaneous breaking $>10^{12}$ GeV which fits very well in our case.
The various ways to solve domain wall problem show that it is possible to overcome it independently. In all the scenarios considered above, it does not prevent the complex VEV of $S$ in generating a complex Yukawa coupling required for CP violation in leptogenesis.
@@
where $m^{(1)}_{\nu}$ and $m^{(2)}_{\nu}$ are the contributions coming from the triplet scalars $\Delta_{1}$ and $\Delta_{2}$, respectively. Here, $u_{a}$ are the VEVs of the neutral components of the triplet scalars $\Delta_{a}$, and it is given by
@@
TBM mixing matrix has gathered considerable attention because of its proximity with the observation of neutrino masses and mixing. However, the absence of Dirac-type CP violating phase asks for possible modifications of the model, which we discuss in the next subsection. Considering the two cases of neutrino mass hierarchy, we can have two sets of constraints on $z_{1}$ and $z_{2}$. From Eq.(2.19), for normal hierarchy (NH), i.e. $m_{3}>m_{2}>m_{1}$, we obtain
@@ 2.2 Parameter space study
$\sin^{2}\theta_{12}$ $0.303^{+0.012}_{-0.012}$ $0.270\rightarrow 0.341$
$\sin^{2}\theta_{23}$ $0.451^{+0.019}_{-0.016}$ $0.408\rightarrow 0.603$
$\sin^{2}\theta_{13}$ $0.02225^{+0.00056}_{-0.00059}$ $0.02052\rightarrow 0.02398$
$\delta/^{\circ}$ $232^{+36}_{-26}$ $144\rightarrow 350$
$\frac{\Delta m^{2}_{21}}{10^{-5}{\rm eV}^{2}}$ $7.41^{+0.21}_{-0.20}$ $6.82\rightarrow 8.03$
$\frac{\Delta m^{2}_{3l}}{10^{-3}{\rm eV}^{2}}$ $+2.507^{+0.026}_{-0.027}$ $+2.427\rightarrow+2.590$
@@
where the phase $\gamma_{1}$ is given in Eq. (2.24). Through the phase, $\gamma_{1}$, it is possible to draw a link between the effective mass and its dependence on the phase $\beta$, which appears in the expressions of neutrino masses as well as in the high-scale CP violation, as it will be elaborated later on. In Fig. (1(b)) we show the dependence of $|m_{ee}|$ on the phase $\beta$. In generating the plot using Eq.(2.36), we have calculated the masses, $m_{1},m_{2}$ and $m_{3}$ using Eq.(2.29), which requires the use of equations (2.27) and (2.28). Also, we ensure the values of $z_{a}$ to be positive as per Eq.(2.20). The plot shows that the allowed effective mass in turn has imposed a bound on the phase $\beta$ $\sim(100^{\circ}-260^{\circ})$. We have used this restricted parameter space of the phase $\beta$ to calculate the flavored CP asymmetries in section (4).
@@ 4.1 Formulation of flavored CP asymmetries
Within the energy range $[10^{9},10^{12}]$ GeV, the flavored CP asymmetries can be expressed as $\epsilon^{\alpha\beta}_{a}(\alpha,\beta=e,\mu)$ and $\epsilon^{\tau\tau}_{a}$, where $a=1,2$ is associated with triplet $\Delta_{1,2}$. Using equations (4.3), (4.4) and (4.6) we calculate, the flavored CP asymmetries. The specific expressions are given in equations (C.8 - C.11) and (C.12 - C.15) in Appendix (C). The dependence of $\beta$ on the CP asymmetries are plotted in the figures (8) and (9) for different values $B^{L}_{1(2)}=0.5,0.05,0.005,0.0005$. The triplet VEV $|u_{1}|$ ($|u_{2}|$) are taken in the range $[0.1-100]$ eV. To estimate the CP asymmetry from the model, we have chosen the small perturbation parameters from the allowed parameter space shown in Fig.(1(a)), as $\varepsilon_{1}\sim-0.125$ and $\varepsilon_{2}\sim 0.075$. We have varied the phase $\beta$ within the range $(100^{\circ}-260^{\circ})$ so that it remains consistent with the latest bound on the effective neutrino mass $|m_{ee}|\lesssim 0.15$ eV. The phase $\sigma_{1}$ is taken in such a way that it satisfies the relations given in Eq.(2.24). The viable values of the parameters $z_{1}$ and $z_{2}$ are calculated as functions of two mass squared differences and phase $\beta$, from the expressions equations (2.27) and (2.28). It is observed that CP asymmetries of the order $\mathcal{O}(10^{-7})$ can be produced for triplet masses $\sim 10^{10}$ GeV. The figures (8) and (9) are generated by taking $M_{1}$ ($M_{2}$) $=2\times 10^{10}$ GeV and taking the mass squared differences in $3\sigma$ range. The scale is suitable for studying two-flavor leptogenesis which we take up in the upcoming sections.
@@ 4.2 Baryogenesis from $\Delta_{1}$ decay
$M_{1}({\rm GeV})$ $B^{L}_{1}$ $\beta(^{\circ})$ $\epsilon^{ee}_{1}/10^{-8}$ $\epsilon^{e\mu}_{1}/10^{-9}$ $\epsilon^{\mu\mu}_{1}/10^{-9}$ $\epsilon^{\tau\tau}_{1}/10^{-9}$ $|\eta_{B}|/10^{-10}$
$0.48$ $178$ $0.70$ $0.84$ $1.01$ $-1.68$ $0.04$
$3\times 10^{10}$ $0.06$ $175$ $2.06$ $2.46$ $2.94$ $-4.91$ $0.29$
$0.005$ $165$ $2.04$ $2.43$ $2.91$ $-4.86$ $0.73$
$0.0005$ $126$ $2.06$ $2.46$ $2.94$ $-4.91$ $1.74$
$0.34$ $178$ $2.79$ $3.32$ $3.98$ $-6.63$ $0.30$
$4\times 10^{10}$ $0.05$ $175$ $2.82$ $3.36$ $4.03$ $-6.71$ $0.44$
$0.005$ $164$ $2.82$ $3.36$ $4.03$ $-6.71$ $1.07$
$0.0006$ $126$ $2.83$ $3.37$ $4.04$ $-6.74$ $2.50$
@@
$M_{1}({\rm GeV})$ $B^{L}_{1}$ $|u_{1}|(\rm eV)$ $\epsilon^{ee}_{1}/10^{-7}$ $\epsilon^{e\mu}_{1}/10^{-8}$ $\epsilon^{\mu\mu}_{1}/10^{-8}$ $\epsilon^{\tau\tau}_{1}/10^{-8}$ $|\eta_{B}|/10^{-10}$
$3\times 10^{10}$ Fig. (4) $0.58$ $4.80$ $4.47$ $5.32$ $6.38$ $-1.06$ $5.29$
@@ 4.3 Baryogenesis from $\Delta_{2}$ decay
$M_{2}({\rm GeV})$ $B^{L}_{2}$ $\beta(^{\circ})$ $\epsilon^{ee}_{2}/10^{-8}$ $\epsilon^{e\mu}_{2}/10^{-9}$ $\epsilon^{\mu\mu}_{2}/10^{-9}$ $\epsilon^{\tau\tau}_{2}/10^{-7}$ $|\eta_{B}|/10^{-10}$
$0.59$ $176$ $-0.66$ $-0.79$ $-0.95$ $0.75$ $0.04$
$1\times 10^{10}$ $0.06$ $172$ $-0.67$ $-0.80$ $-0.96$ $0.76$ $0.07$
$0.005$ $157$ $-0.67$ $-0.80$ $-0.96$ $0.75$ $0.86$
$0.0006$ $126$ $-0.73$ $-0.87$ $-1.05$ $0.75$ $3.26$
$0.53$ $177$ $-1.29$ $-1.53$ $-1.84$ $1.45$ $0.01$
$2\times 10^{10}$ $0.05$ $176$ $-1.42$ $-1.68$ $-2.02$ $1.59$ $0.03$
$0.006$ $156$ $-1.42$ $-1.69$ $-2.03$ $1.58$ $1.66$
$0.0005$ $126$ $-1.42$ $-1.69$ $-2.03$ $1.45$ $7.51$
@@
$M_{2}({\rm GeV})$ $B^{L}_{2}$ $|u_{2}|(\rm eV)$ $\epsilon^{ee}_{2}/10^{-8}$ $\epsilon^{e\mu}_{2}/10^{-9}$ $\epsilon^{\mu\mu}_{2}/10^{-9}$ $\epsilon^{\tau\tau}_{2}/10^{-7}$ $|\eta_{B}|/10^{-10}$
$1.4\times 10^{10}$ Fig. (5) $0.0006$ $48.1$ $-1.00$ $-1.20$ $-1.44$ $1.04$ $4.92$